OpenStudy (anonymous):

1+ tan^2 /1+cos^2= tan^2 prove the identity

OpenStudy (mertsj):

\[\frac{1+\tan ^2}{1+\cos ^2}=\tan ^2\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

im proving for the identity

OpenStudy (mertsj):

\[1+\frac{\tan^2}{1+\cos^2}=\tan^2\]

OpenStudy (mertsj):

Which is it?

OpenStudy (anonymous):

the csc - sin (theta)

OpenStudy (mertsj):

I would love to help you but since you refuse to tell me what the problem is, I have no way of knowing which problem to work.

OpenStudy (anonymous):

:( thanks